OpenStudy (anonymous):

SUPER LOST! can some please help me with my geometry problems! -> A lighthouse built at sea level is 150 feet high. The angle from the top of the lighthouse to the buoy is 65 degrees. What is the distance from the buoy to the foot of the lighthouse? & What is the distance from the buoy to the top of the lighthouse?

OpenStudy (anonymous):

x= 150 / sin 25

OpenStudy (jagatuba):

There are three equations that you must remember when solving right triangles: sin A=o/h cos A=a/h tan A=o/a Where A is the angle, o is the opposite leg, and a is the adjacent leg. With these three equations you can solve any right triangle given any two of the other factors.

OpenStudy (jagatuba):

Oops, I left out h is the hypotenuse. lol sorry.

OpenStudy (dumbcow):

 tan65 = x/150, solve for x cos65 = 150/h, solve for h